OpenStudy (anonymous):

55.27 is the same as 55 ____' 12".

OpenStudy (psymon):

It may not be needed, but thought I would add on. 1 minute is (1/60) of a degree and 1 minute is (1/3600) of a degree. To find out what .27 is in minutes and degrees, you need to have 60 divisions of .27. .27 can be rewritten as (27/100). Divided by 60 is how we come across \[\frac{ 27 }{ 100 }(60)=\frac{ 81 }{ 5 }=16\frac{ 1 }{ 5 }\] The 16 is how many minutes you have. If you were not given the seconds, we would have to say that we have a remainder of (1/5) of a minute. Since 1 minute is (1/60), one-fifth of that is (1/300). Now we would need to see how many seconds that gives us. 1 second is (1/3600) of a degree, so we would need to divide (1/300) by (1/3600). Doing this gives us: \[\frac{ 1 }{ 300 }\div \frac{ 1 }{ 3600 }=\frac{ 1 }{ 300 }(3600)=12\]Which is how we got the 12 seconds part.